# Dynamic Attention-Guided Diffusion for Image Super-Resolution

### v3

@@ 4 Experiments
We start by analyzing different methods for obtaining attention maps for YODA. Then, we evaluate YODA and compare its performance in tandem with SR3 [38] for face-only, and SRDiff [25] for general SR. We chose SR3 and SRDiff because they are the most prominent representative diffusion models for image SR in the respective tasks, where YODA can be integrated straightforwardly. However, YODA can be theoretically applied to any existing method. We present quantitative and qualitative results for both tasks. Our method achieves high-quality results for both tasks and outperforms the baselines using standard metrics such as PSNR, SSIM, and LPIPS [34]. All experiments were run on a single NVIDIA A100-80GB GPU. In the supplementary materials, we discuss the complexity of YODA and explore its potential synergies with other diffusion model research.
Figure 2: Comparison of attention maps (blue = low attention; yellow = high attention). Top row denotes maps derived from ResNet-50 using DINO. It shows various attention head outputs and the max aggregation of all attention maps (MAX). Bottom row denotes non-learnable methods, namely Gaussian, Edge-based, and using SIFT’s points of interest. The maps are subsequently used to produce time-dependent binary masks.
4.1 Analysis of Attention Maps
Since the dynamic masking in YODA relies on attention maps, we thoroughly evaluate different choices. We considered the pre-trained attention heads from the last layer of DINO with the respective neural network, i.e., ResNet and ViT [20, 14]. For ResNet-50, we used a dedicated method to extract the attention maps from its weights [19]. A qualitative visual comparison of attention maps generated with DINO and ResNet-50 is shown in Figure 2. It shows that YODA captures and highlights perceptually essential areas. More visual results can be found in the supplementary materials. In addition to different neural networks, we test non-learnable methods to extract attention maps, also shown in Figure 2:
  • Edge-based: Using the Canny edge detector, the attention maps are defined by the edges of the image, where adjacent and near edges are connected and blurred.
Table 1: Results of using different attention maps with SR3+YODA for $16\times 16\rightarrow 128\times 128$ on CelebA-HQ. Aggregating the attention maps extracted with DINO and ResNet-50 backbone under the MAX strategy performs best.
Attention Maps for YODA PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
PULSE (No Mask) 16.88 0.440 n.a.
FSRGAN (No Mask) 23.01 0.620 n.a.
SR3 Reported (No Mask) 23.04 0.650 n.a.
SR3 Reproduced (No Mask) 22.35 0.646 0.082
Non-learnable, fixed: Gaussian 22.13 0.602 0.260
Non-learnable, adaptive: Edge-based 22.93 0.648 0.151
Non-learnable, adaptive: SIFT 22.84 0.678 0.095
ViT-S/8 Attention-Head 0 22.91 0.650 0.105
@@ 4.1 Analysis of Attention Maps
ResNet-50 Attention-Head 0 22.82 0.649 0.115
@@
Table 1 presents the results of our study with several baselines and masking variants for $16\times 16\rightarrow 128\times 128$ scaling on the CelebA-HQ dataset. The straightforward Gaussian approach performs worst as it does not adapt to image features. The edge-based segmentation and SIFT methods showed improved performance over the reproduced baseline using a small batch size. However, they underperformed relative to the reported SR3 [38] results, which used a larger batch size. In contrast, using DINO [6] to extract attention maps improves performance, regardless of the choice of backbone (ResNet-50 and ViT-S/8). We tested individual attention heads (0 to 5) independently, along with combination strategies that include averaging (AVG) and selecting the maximum value (MAX). The MAX combination achieved the best results compared to individual heads or the AVG combination. DINO [6] with a ResNet-50 backbone performs best, and we used the MAX aggregation strategy for all main experiments.
@@
The left part of Figure 3 provides more information on the ratio of diffused pixels using our time-dependent masking. The upper bound is 100%, which is the case if diffusion is applied across all pixel locations throughout every time step (as in standard diffusion). Thus, any result under 100% shows that not all pixels are diffused during all time steps. As can be seen, DINO [6] coupled with ResNet-50 produces higher attention values, thereby leading to more total pixel updates. Also, the high variance of the ResNet-50 backbone indicates a high adaption capability. It can employ 100% of the updates for some samples, a characteristic not observed with the ViT-S/8 backbone. Nevertheless, the improved performance with the ViT-S/8 backbone compared to non-learnable methods and its low ratio of diffused pixels makes it an attractive candidate for future work regarding inference speed optimization based on sparser diffusion.
The right part of Figure 3 shows the ratio of diffused pixels depending on time steps with the MAX aggregation for the ResNet-50 and the ViT-S/8 backbone. As the backward diffusion goes from $T$ to 00, ResNet-50 backbone initiates and incorporates the refinement of the whole image areas more quickly than ViT-S/8. For the first 200 time steps, less than 20% of the image area is addressed by the attention map derived by ViT-S/8, whereas ResNet-50 has already developed to 100%. Therefore, this observation leads to our assumption that the superior performance of the ResNet-50 backbone may be attributed to its faster progression in refining the image. In other words, it performs better by progressing from 0% to 100% of the image area at an earlier stage. Intermediate diffusion results can be found in the supplementary materials.
Table 2: Results on face SR with 4× scaling ($64\times 64\rightarrow 256\times 256$) and 8× scaling ($64\times 64\rightarrow 512\times 512$) on CelebA-HQ. The models were trained for 1M steps on FFHQ and a reduced batch size of 4 and 8 instead of 256 to fit on a single A100-80GB GPU.
Scaling Model PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
SR3 [38] 17.98 0.607 0.138
SR3 + YODA 26.33 0.838 0.090
PartDiff ($K$=25) [48] - - 0.222
PartDiff ($K$=50) [48] - - 0.217
SR3 [38] 17.44 0.631 0.147
SR3 + YODA 25.04 0.800 0.126
@@ 4.2 Face Super-Resolution
4.2.1 Details
We use the FFHQ dataset [23] for training, which contains 50,000 high-quality facial images. We adopted the AdamW [29] optimizer, using a weight decay of 0.0001 and a learning rate of 5e-5. The number of sampling steps is set to $T_{\text{train}}=500$. We use the CelebA-HQ dataset [22] for evaluation, which contains 30,000 facial images. The number of sampling steps is set to $T_{\text{eval}}=200$. We trained all models for 1M iterations as in SR3 [38]. We evaluated three scenarios: $16\times 16\rightarrow 128\times 128$, $64\times 64\rightarrow 256\times 256$, and $64\times 64\rightarrow 512\times 512$. Due to hardware limitations and missing quantitative results in the original publication of SR3, our experiments required a reduction from the originally used batch size of 256: we used a batch size of 4 for the $64\times 64\rightarrow 512\times 512$, and 8 for the $64\times 64\rightarrow 256\times 256$ scenario to fit on a single A100-80GB GPU.
4.2.2 Results
@@
The results in Table 2 show a significant improvement when SR3 is coupled with YODA across all examined metrics. We explain the poor performance of SR3 with a phenomenon that is also observed by other authors [42, 8]. SR3 suffers from a color shift issue, which we attribute to the reduced batch size necessitated by our hardware limitations. Examples are shown in Figure 4 and Figure 5. This color shift manifests in a pronounced deviation in pixel-based metrics such as PSNR and SSIM but only slightly decreased perceptual quality in terms of LPIPS. Our results suggest that YODA’s role extends beyond mere performance enhancement. It actively mitigates the color shift issue and stabilizes training, especially when faced with hardware constraints. With YODA, SR3 can be trained with a much smaller batch size and still achieves strong performance. Figure 5 offers another qualitative comparison between SR3 and SR3+YODA, highlighting subtle yet important differences, e.g., around the eyes, mouth, and hair.
@@
$\uparrow$ PSNR: 11.811
$\uparrow$ SSIM:   0.3311
$\downarrow$ LPIPS: 0.2847
@@
$\uparrow$ PSNR: 18.069
$\uparrow$ SSIM:   0.6588
$\downarrow$ LPIPS: 0.2628
@@
Figure 4: A comparison of LR, HR, SR3, and SR3+YODA images in the $64\times 64\rightarrow 256\times 256$ setting (4x). SR3 suffers from color shift issues, as also observed by [42, 8]. Our method solves this issue and produces higher quality reconstructions.
LR
SR3
$\uparrow$ PSNR: 23.061
$\uparrow$ SSIM:   0.6208
$\downarrow$ LPIPS: 0.0504
SR3+YODA
$\uparrow$ PSNR: 23.289
$\uparrow$ SSIM:   0.6334
$\downarrow$ LPIPS: 0.0502
HR
Figure 5: A comparison of LR, HR, SR3, and SR3+YODA images in the $16\times 16\rightarrow 128\times 128$ setting (8x). The color shift in SR3 can still be observed (e.g., see upper right corners). YODA produces higher-quality images without color shift issues.
@@ 4.3 General Super-Resolution
4.3.1 Details
The experimental design follows SRDiff [25] and its hyperparameters, which are originally based on the experimental design of SRFlow [32]. For training, we employed 800 2K resolution HQ images from DIV2K [1] and 2,650 images from Flickr2K [41]. For testing, we used the DIV2K validation set of 100 images. Furthermore, we evaluated SR3, which was not originally tested on DIV2K. As in SRDiff, we extracted $40\times 40$ sub-images with a batch size of 16, AdamW [29], a channel size of 64 with channel multipliers [1, 2, 2, 4] and $T=100$.
Figure 6: Example zoom-in regions of images (first row) from the DIV2K dataset. LPIPS is reported in the boxes (the lower, the better). YODA consistently produces better texture and more high-frequency details.
Table 3: Quantitative results of 4× general SR on the DIV2K validation set.
@@ 4.3.1 Details
4.3.2 Results
Table 3 shows the 4× scaling general image SR results on the DIV2K validation set. The reported values include regression-based methods, which typically yield higher pixel-based scores than generative approaches [38]. This disparity is due to PSNR/SSIM penalizing misaligned high-frequency details, a known and significant challenge in the wider SR field [34].
We observe that results from SR3 underperform compared to YODA. Equipping SR3 with YODA improves image quality even without an extensive hyperparameter search. We hypothesize that a vanilla SR3 strongly depends on larger batch sizes and longer diffusion times.
Combining SRDiff with YODA demonstrates improved performance in PSNR (+0.21db) and SSIM (+0.01), with a minor deterioration in LPIPS (+0.01). Thus, our approach excels in pixel-based metrics but sees a marginal decline in the perceptual metric. Nonetheless, when looking qualitatively at the predictions, one can observe that SRDiff benefits from YODA, as shown in Figure 6. YODA produces much better LPIPS values for perceptually essential areas, i.e., hair and cars, but falls short in background areas, i.e., blurry grass or dark ground. More results can be found in the supplementary materials.
Overall, YODA’s strengths are more significant when combined with SR3 than with SRDiff. A critical distinction between SR3 and SRDiff is the handling of conditional information, i.e., the LR image, which we identify as a potential contributor to the reduced perceptual score LPIPS. SRDiff employs an LR encoder that generates an embedding during the denoising phase. Meanwhile, SR3 directly uses the LR image during the backward diffusion.
Furthermore, SR3 operates on the full LR images, whereas SRDiff diffuses residuals. Hence, the attention maps generated beforehand might not accurately capture the essential regions of the sparse residual inputs. Another possible reason for the lower perceptual score could be the image size during DINO training, i.e., $224\times 224$ for the teacher network, which is much smaller than the 2K resolution images in DIV2K. As such, fine-tuning DINO on larger-scale images might be essential to capture more meaningful semantic features.
@@ 5 Conclusion
In this work, we presented “You Only Diffuse Areas" (YODA), an attention-guided diffusion-based image SR approach that emphasizes specific areas through time-dependent masking. YODA first extracts attention maps that reflect the pixel-wise importance of each scene using a self-supervised, general-purpose vision encoder. The attention maps are then used to guide the diffusion process by focusing on key regions in each time step while providing a fusion technique to ensure that masked and non-masked image regions are correctly connected between two successive time steps. This targeting allows for a more efficient transition to high-resolution outputs, prioritizing areas that gain the most from iterative refinements, such as detail-intensive regions. Beyond performance enhancement, YODA stabilizes training and mitigates the color shift issue that emerges when a reduced batch size constrains vanilla SR3. As a result, YODA consistently outperforms strong SR baselines such as SR3 and SRDiff, while requiring less computational resources by using smaller batch sizes.

### v4

@@ 4 Experiments
We start by analyzing different methods for obtaining attention maps for YODA. Then, we evaluate YODA’s performance in tandem with SR3 [46] and DiffBIR [31] for face, as well as SRDiff [28] for general SR. We chose SR3, DiffBIR, and SRDiff because they are the most prominent representative diffusion models for image SR in the respective tasks, where YODA can be integrated straightforwardly. However, YODA can be theoretically applied to any existing method. We present quantitative and qualitative results for both tasks, demonstrating YODA’s high-quality results compared to the baselines using standard metrics such as PSNR, SSIM, and LPIPS [40]. All experiments were run on a single NVIDIA A100-80GB GPU. In the supplementary materials, we discuss the complexity of YODA and explore its potential synergies with other diffusion models. Also, we used a lower bound hyperparameter (see Section 3.2) of $l=0.2$ in all experiments and were inspired by the rate-distortion trade-off presented by Ho et al. [23] that reaches the semantic compression stage at roughly $t=T-0.2\cdot T$.
Attention Maps PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
no masks PULSE 16.88 0.440 n.a.
FSRGAN 23.01 0.620 n.a.
SR3 Reported 23.04 0.650 n.a.
SR3 Reproduced 22.35 0.646 0.082
innate Gaussian 22.13 0.602 0.260
Edge-based 22.93 0.648 0.151
SIFT 22.84 0.678 0.095
DINO with ViT-S/8 Attention-Head 0 22.91 0.650 0.105
@@
DINO with ResNet-50 Attention-Head 0 22.82 0.649 0.115
@@
Figure 2: (Left) Comparison of various methods to extract attention maps used for our method (blue = low attention; yellow = high attention). Top row denotes maps derived from ResNet-50 using DINO. It shows various attention head outputs and the max aggregation of all attention maps (MAX). Bottom row denotes non-learnable methods, namely Gaussian, Edge-based, and using SIFT’s points of interest. (Right) Comparison of different attention maps with SR3+YODA for $16\rightarrow 128$ on CelebA-HQ. Aggregating the attention maps extracted with DINO and ResNet-50 backbone under the MAX strategy performs best. The attention maps are then used for dynamic binary masking.
4.1 Choosing Good Attention Maps
YODA relies on attention maps. Thus, we thoroughly evaluated different choices. We considered the pre-trained attention heads from the last layer of DINO with the respective backbone model, i.e., ResNet [22] and ViT [16]. For ResNet-50, we used a dedicated method to extract the attention maps from its weights [21]. A qualitative comparison of attention maps generated with DINO and ResNet-50 is shown in Figure 2 (left), demonstrating that YODA highlights perceptually essential areas (more visual results are in the appendix). Additionally, we test non-learnable methods to extract attention maps, also shown in Figure 2 (left):
  • Edge-based: Using the Canny edge detector, the attention maps are defined by the edges of the image, where close edges are connected and blurred.
DINO masks perform best: Figure 2 (right) presents our study on several baselines and masking variants. The straightforward Gaussian approach performs worst as it does not adapt to image features. The edge-based segmentation and SIFT methods improve the performance over the reproduced baseline using a small batch size. However, they underperform relative to the reported SR3 results, which used a larger batch size. In comparison, using DINO’s attention maps for YODA shows significant improvements. We tested individual attention heads (0 to 5) independently, along with combination strategies that include averaging (AVG) and selecting the maximum value (MAX). The MAX combination achieved the best results compared to individual heads or the AVG combination.
@@ 4.1 Choosing Good Attention Maps
ResNet produces more sensitive masks with more pixel updates: Figure 3 (left) investigates the ratio of diffused pixels using our time-dependent masking. The upper bound is 100%, where diffusion is applied across all locations in every time step (standard diffusion). Any result under 100% shows that not all pixels are diffused during all time steps. As can be seen, DINO with ResNet-50 produces higher attention values, resulting in more total pixel updates. Also, the high variance indicates a high adaptability. It can employ 100% for some samples, a characteristic not observed with ViT-S/8. Nevertheless, the ViT-S/8’s improved performance compared to non-learnable methods and its low ratio make it attractive for future work on optimized inference speed based on sparser diffusion, e.g., LazyDiffusion [42].
ResNet progresses faster towards the whole image: Figure 3 (right) shows the ratio of diffused pixels depending on time steps with the MAX aggregation for ResNet-50 and ViT-S/8. As the backward diffusion progresses from $T$ to 00, ResNet-50 initiates and incorporates the refinement of the whole image areas more quickly than ViT-S/8. For the first 200 time steps, the attention map derived by ViT-S/8 addresses less than 20% of the image area, whereas ResNet-50 has already developed to 100%. Therefore, we assume that the ResNet-50 backbone’s superior performance is attributed to its faster progression toward refining the whole image. Intermediate diffusion results and error maps can be found in the supplementary materials.
Summary: As DINO with ResNet-50 and MAX aggregation performs best, we used it for all remaining main experiments.
@@ 4.2 Face Super-Resolution
We use FFHQ [25] for training and CelebA-HQ for testing [24]. All SR3 models were trained for 1M iterations as in Saharia et al. [46]. We evaluated three scenarios with bicubic degradation: $16\rightarrow 128$, $64\rightarrow 256$, and $64\rightarrow 512$. Due to hardware limitations and missing quantitative results in the original publication of SR3, our experiments required a reduction from the originally used batch size of 256: we used a batch size of 4 for the $64\rightarrow 512$, and 8 for the $64\rightarrow 256$ scenario to fit on a single A100-80GB GPU. For blind face SR (unknown degradation between LR and HR, $64\rightarrow 256$), we follow Lin et al. and test YODA with DiffBIR [31], which uses more complex degradation models (e.g., blurring), like introduced by Real-ESRGAN and others [54, 58].
Scaling Type Model PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$
Regression RRDB [54] 27.77 0.870 0.151 67.46
Diffusion SR3 [46] 17.98 0.607 0.138 80.72
Diffusion SR3 + YODA 26.33 0.838 0.090 59.99
Regression RRDB [54] 26.91 0.780 0.220 62.85
Diffusion PartDiff ($K$=25) [62] n.a. n.a. 0.222 n.a.
Diffusion PartDiff ($K$=50) [62] n.a. n.a. 0.217 n.a.
Diffusion SR3 [46] 17.44 0.631 0.147 66.20
Diffusion SR3 + YODA 25.04 0.800 0.126 53.95
Diffusion DiffBIR [31] 24.49 0.717 0.247 115.22
(blind) Diffusion DiffBIR + YODA 24.56 0.718 0.245 111.93
Table 1: Face SR results with 4× scaling ($64\rightarrow 256$) and 8× scaling ($64\rightarrow 512$) on CelebA-HQ with SR3 (non-blind) and DiffBIR (blind means unknown degradation type) standalone and combined with YODA. Note that RRDB [54] is also reported and that regression-based methods typically yield higher pixel-based scores (PSNR/SSIM) than generative approaches [46]. SR3 was trained for 1M steps and a reduced batch size of 4 and 8 instead of 256 to fit on a single A100-80GB GPU. Note that DiffBIR uses diffusion after an initial, regression-based predictor with only 50 sampling steps. Thus, smaller relative improvements are expected.
Results: Table 1 shows significant improvements when SR3 (face SR) or DiffBIR (blind face SR) is coupled with YODA across all metrics. DiffBIR applies a diffusion process following an initial regression-based predictor and uses only 50 sampling steps. Consequently, smaller relative improvements are expected, but it shows that YODA is also efficient for more complex degradation models.
@@
We explain the poor performance of SR3 with a phenomenon also observed by other works [50, 9]: color shifting, which we attribute to the reduced batch size necessitated by limited hardware access. Fitting SR3 on a single A100-80GB GPU required reducing the batch size from 256 to 8. Examples are shown in Figure 4 and in the appendix. Color shifting manifests in pronounced deviation in pixel-based metrics (PSNR/SSIM) but only slightly decreased perceptual quality (LPIPS). Our results suggest that YODA’s role extends beyond mere performance enhancement. It actively mitigates color shifting and stabilizes training, especially when faced with hardware constraints. Due to selective refinement, YODA maintains more of the LR image in less complex areas, reducing the risk of introducing artifacts/shifts during denoising while enhancing details in complex regions. With YODA, SR3 can be trained with a much smaller batch size.
Results without color-shifting: Table 2 shows the same experiments but with the channel-wise mean normalized by the ground-truth data, thus disentangling image quality improvements from color bias. While this procedure is not feasible in practice, it highlights YODA’s significantly improved performance beyond reducing color shifting.
Analysis across attention regions: In Figure 5, we analyze the LPIPS scores across different attention value intervals within a single attention map $\mathbf{A}$ (using the MAX aggregation, 0.01 interval size). These intervals represent varying attention levels assigned by DINO to different regions of the image. The high LPIPS scores associated with bicubic upsampling in high-attention areas underscore the effectiveness of DINO in capturing perceptually significant regions. As a result, YODA significantly improves LPIPS scores across all attention regions, particularly in regions with higher attention values (highlighted by the trend curve). Please refer to the appendix for more details.
User study: In addition to the quantitative results on $16\rightarrow 128$ (8×) provided in Figure 2 and inspired by Saharia et al., we conducted a user study. We selected 50 random test images, asked 45 participants which SR prediction is preferred, and plotted the preferences per image (see appendix). As a result, YODA was preferred 55.7% of the time over SR3 (44.3%). More details are in the appendix.
@@
$\uparrow$ PSNR:
$\uparrow$ SSIM:
$\downarrow$ LPIPS:
@@
11.811
0.3311
0.2847
@@
18.069
0.6588
0.2628
@@
Figure 4: SR3 and SR3+YODA reconstructions, $64\rightarrow 256$ (4×). SR3 suffers from color shifting, as also observed by [50, 9]. YODA solves this issue and produces higher-quality reconstructions.
Scaling Model PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$
SR3 [46] 28.3 0.67 0.10 74.99
SR3 + YODA 31.6 0.87 0.05 58.34
SR3 [46] 29.3 0.71 0.10 55.94
SR3 + YODA 31.5 0.84 0.07 49.02
Table 2: Results without color-shifting on face SR by normalizing the channel-wise means to those of the HR ground-truth data for 4× scaling ($64\rightarrow 256$) and 8× scaling ($64\rightarrow 512$).
@@ 4.3 General Super-Resolution
The experimental setup follows SRDiff [28], based on the setup of SRFlow and bicubic degradation [36]. For training, we employed 800 2K resolution images from DIV2K [1] and 2,650 from Flickr2K [49]. For testing, we used the DIV2K validation set (100 images). Moreover, we evaluated SR3, which was not originally tested on DIV2K.
Figure 5: Regional LPIPS comparison across normalized attention values for CelebA, $64\rightarrow 256$ (4×). We use 0.01 intervals and fit a polynomial through the means. High-attention areas are perceptually relevant and correspond to more difficult pixels (higher LPIPS). YODA reaches better scores, especially within high-attention areas. Note that dynamic masking stops around $t\approx 0.6\cdot T$, see Figure 3.
Results: Table 3 shows the 4× general SR results on the DIV2K val. The reported values include regression-based methods, which typically yield higher pixel-based scores than generative models [46]. The disparity is due to PSNR/SSIM penalizing misaligned high-frequency details, a known challenge in the wider SR field [40]. We observe that results from SR3 underperform without YODA. We hypothesize that SR3 strongly depends on larger batch sizes and longer diffusion times. Combining SRDiff with YODA improves the quality in PSNR (+0.21db) and SSIM (+0.01), with a minor deterioration in LPIPS (+0.01). Nonetheless, when looking qualitatively at the predictions, one can observe that SRDiff strongly benefits from YODA, as shown in Figure 6. YODA produces much better LPIPS values for perceptually essential areas, i.e., hair and cars, but falls short in background areas, i.e., blurry grass or dark ground. The appendix contains more visual results.
Discussion: Overall, YODA’s strengths are more significant when combined with SR3 than with SRDiff. A critical distinction between SR3 and SRDiff is the handling of conditional information, i.e., the LR image, which we identify as a potential contributor to the reduced perceptual score LPIPS. SRDiff employs an LR encoder that generates an embedding during the denoising phase. Meanwhile, SR3 directly uses the LR image during the backward diffusion.
@@
Table 3: Quantitative results of 4× general SR on DIV2K val. YODA improves across pixel-centric metrics such as PSNR and SSIM, but yields a marginal decline in LPIPS.
@@ 5 Conclusion
In this work, we presented “You Only Diffuse Areas” (YODA), an attention-guided diffusion-based image SR approach that emphasizes essential areas through time-dependent masking. YODA first extracts attention maps that reflect the pixel-wise saliency of each scene using a self-supervised, general-purpose vision encoder. The attention maps are then used to guide the diffusion process by focusing on key regions in each time step while providing a fusion technique to ensure that masked and non-masked image regions are correctly connected between two successive time steps. This targeting allows for a more efficient transition to high-resolution outputs, prioritizing areas that gain the most from iterative refinements, such as detail-intensive regions. Beyond better performance, YODA stabilizes training and mitigates the color shift issue when a reduced batch size constraints the underlying diffusion model. As a result, YODA consistently outperforms strong baselines like SR3, DiffBIR, and SRDiff while requiring less computational resources by using smaller batch sizes.
Figure 6: Zoom-in regions of DIV2K images (first row). LPIPS is reported in the boxes (the lower, the better). YODA consistently produces better texture and more high-frequency details.